OpenStudy (anonymous):

An airplane leaves an airport and flies due west 150 miles and then 170 miles in the direction S 49.17°W. How far is the plane from the airport (round to the nearest mile)?

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

use the law of cosines to find the value of c  c^2 = a^2 + b^2 - 2*a*b*cos(C) c^2 = 150^2 + 170^2 - 2*150*170*cos(49.17) .... .... c = ???

jimthompson5910 (jim_thompson5910):

oh wait, C isn't 49.17 C is 49.17+90 = 139.17 so it should be c^2 = 150^2 + 170^2 - 2*150*170*cos(139.17)